OpenStudy (anonymous):

Graph the solution set of |-4w+2|=14 Show work please

OpenStudy (anonymous):

|-4w+2|=14 2|-w+1|=14 |-w+1|=7 -w+1=7 or -w+1=-7 -w=6 or -w=-8 w=-6 or w=8

OpenStudy (anonymous):

thanks. What confuses me, is I thought absolute values had to be positive, and I wasn't sure how to handle the -4w. From your answer, it looks like I solve it like a regular expression. Is this correct?

OpenStudy (anonymous):

There is an error in the solution. When you factored out a 2 from the absolute value sign, what remains inside the absolute value is -2w+1, not -w+1. Solutions will change.

OpenStudy (anonymous):

@jz55

OpenStudy (anonymous):

@WHAT?!

OpenStudy (anonymous):

Thank you for catching this.

OpenStudy (anonymous):

Welcome.